OpenStudy (anonymous):

Determine whether each of the following trinomials is a perfect square. x2-14x +49

OpenStudy (experimentx):

(x-7)^2

OpenStudy (hoblos):

x2-14x +49 = x2 -2(7)(x) + 7^2 = (x-7)^2

OpenStudy (anonymous):

that would be an x squared. I couldn't figure out how to make the 2 smaller.

OpenStudy (hoblos):

you can use x^2

OpenStudy (anonymous):

how did you find (x-7)(x+7)?

OpenStudy (anonymous):

Pardon me... I have forgotten how to do this. Bleh!

OpenStudy (anonymous):

The patterns you are looking for is \[(x+a)^2=x^2+2ax+a^2\]or\[(x-a)^2=x^2-2ax+a^2\]

OpenStudy (hoblos):

well (x-7)(x+7) is wrong because (x-7)(x+7)= x^2 - 7^2 which is different that the given!!

OpenStudy (anonymous):

okay, so it's (x-7)(x-7) correct?

OpenStudy (ajprincess):

yes

OpenStudy (anonymous):

can somebody please show me how to get this answer... I am still not sure how we found it.

OpenStudy (ajprincess):